# Accurate and precise quantum computation of valence two-neutron systems

### v1

@@ I Introduction
There are several pioneering applications of quantum algorithms to nuclear many-body problems, mostly on the Lipkin model [15, 16, 17] and on shell-model Hamiltonians [18, 19, 20]. However, this is still developing, and further research is needed to usher in the era of quantum computers for a more general class of nuclear many-body problems.
This study explores how to solve the valence two-neutron systems accurately and precisely with the present NISQ devices. By taking into account the symmetries of the system in an explicit manner, and selecting appropriate ansatz (circuits), the optimization method, and the error mitigation technique, we will show that the proposed method is proven to give very accurate results of the ground-state energies with fewer CNOT gates than the previous works utilizing e.g., Unitary Coupled Cluster (UCC) ansatz and its variants [18, 19, 20].
@@ II.1 Shell-model calculation
The systems of interest are the ground states of two-neutron systems within the $p$, $sd$, and $pf$ shell, i.e., ${}^{6}\mathrm{He}$, ${}^{18}\mathrm{O}$, and ${}^{42}\mathrm{Ca}$, respectively. Table 1 summarizes the exact results for these nuclei with phenomenological effective interactions, Cohen-Kurath ($p$ shell) [24], USDB ($sd$ shell) [25], and GXPF1A ($pf$ shell) [26]. One can obtain the corresponding interaction files and reproduce these results with shell-model codes such as KSHELL [27, *KSHELL2] and NuclearToolkit.jl [29, *Repo_NuclearToolkit.jl]. For these nuclei, the calculation with a classical computer is easy, but it is still a good test ground for the quantum computer and error mitigation techniques.
@@ III.1 Measurements of pair occupation numbers
From noise-free results (sim. FTQC) in these figures, one can see that the expectation values of the states occupying the same jj-coupled orbital are identical within a statistical error. As a whole, ground-state wave functions of the nuclei of interest concentrate on the configurations occupying the lowest jj-coupled orbital, which is usually to be the orbital with the highest angular momentum $j$ among the adopted valence orbitals, i.e., $0d_{5/2}$ ($0f_{7/2}$) for the $sd$ ($pf$) shell. Due to the errors of the quantum devices or simulator mimicking them, the sim. NISQ/Real results are slightly different from the noise-free ones. It should be noted that sim. NISQ and Real results may vary depending on the quantum device employed and its calibration status.
@@ V Summary and Outlook
The target systems discussed in this work are two-neutron systems in the valence space, which are rather simple. The extension of this work to the systems with more valence nucleons is important, and is left for future work. In addition, it is also important to develop other error mitigation techniques, which are more general and applicable to various systems.
The “pair-wise” form has been used in quantum chemistry referred to as the “antisymmetrized geminal power (AGP)” to be used as a reference state of the coupled cluster calculation for the strongly correlated systems to describe two-particle correlation more properly than a single Slater determinant [36, 37, 38, 39, 40, 32, 31], where a geminal is a two-body correspondence of a one-body orbital. The operator of the geminal is quite similar to the Bardeen-Cooper-Schrieffer (BCS) one and, actually, it is pointed out in Ref. [39] that the AGP wave function is strongly related to the particle-number conserving BCS wave function [41]. Therefore, it is natural to use this AGP ansatz for even-even nuclei as a starting point since the valence nucleons of even-even nuclei form Cooper pairs [42]. It should be noted that the ground-state wave function of the BCS state can also be written by a Pfaffian, instead of as Slater determinant [43, 44, 45]; a Pfaffian wave function and an AGP one are also based on the same philosophy. Hence it is expected that the AGP like ansatz is also useful for even-even nuclei.
@@ Acknowledgements.
We acknowledge the use of IBM Quantum services for this work. T.N. thanks Osamu Sugino for the fruitful discussion on AGP. This research was supported in part by JSPS Grant-in-Aid for Scientific Research (Grant Nos. JP19H00869, JP20H05670, JP21K18903, JP22H05025, JP22K14030, JP22K20372, JP23H01845, JP23H04526, JP23K01845, JP23K03426, JP24K17057), JST COI-NEXT (Grant No. JPMJPF2221), and MEXT Q-LEAP (Grant No. JPMXS0118067246). S.Y. acknowledges JGC-Saneyoshi Scholarship Foundation. T.N. acknowledges the RIKEN Special Postdoctoral Researcher Program.

### v2

@@ I Introduction
There are several pioneering applications of quantum algorithms to nuclear many-body problems, mostly on the Lipkin model [15, 16, 17] and on shell-model Hamiltonians [18, 19, 20, 21, 22]. However, this is still developing, and further research is needed to usher in the era of quantum computers for a more general class of nuclear many-body problems.
This study explores how to solve the valence two-neutron systems accurately and precisely with the present NISQ devices. By taking into account the symmetries of the system in an explicit manner, and selecting appropriate ansatz (circuits), the optimization method, and the error mitigation technique, we will show that the proposed method is proven to give very accurate results of the ground-state energies with fewer CNOT gates than the previous works utilizing e.g., Unitary Coupled Cluster (UCC) ansatz and its variants [18, 19, 20, 21, 22].
@@ II.1 Shell-model calculation
The systems of interest are the ground states of two-neutron systems within the $p$, $sd$, and $pf$ shell, i.e., ${}^{6}\mathrm{He}$, ${}^{18}\mathrm{O}$, and ${}^{42}\mathrm{Ca}$, respectively. Table 1 summarizes the exact results for these nuclei with phenomenological effective interactions, Cohen-Kurath ($p$ shell) [26], USDB ($sd$ shell) [27], and GXPF1A ($pf$ shell) [28]. One can obtain the corresponding interaction files and reproduce these results with shell-model codes such as KSHELL [29, *KSHELL2] and NuclearToolkit.jl [31, *Repo_NuclearToolkit.jl]. For these nuclei, the calculation with a classical computer is easy, but it is still a good test ground for the quantum computer and error mitigation techniques.
@@ III.1 Measurements of pair occupation numbers
From noise-free results (sim. FTQC) in these figures, one can see that the expectation values of the states occupying the same jj-coupled orbital are identical within a statistical error. This sim. FTQC results is by construction identical to the exact results. As a whole, ground-state wave functions of the nuclei of interest concentrate on the configurations occupying the lowest jj-coupled orbital, which is usually to be the orbital with the highest angular momentum $j$ among the adopted valence orbitals, i.e., $0d_{5/2}$ ($0f_{7/2}$) for the $sd$ ($pf$) shell.
On the other hand, the state preparation on the noisy simulators (sim. NISQ) and the real device (Real) show certain deviations from the exact results. It should be noted that sim. NISQ and Real results may vary depending on the quantum device employed and its calibration status.
@@
In the Table 2, we summarize the observed distribution of the particle numbers in the results on the ibm_brisbane device. The $\bar{N}_{\scriptstyle{\text{{{{occ}}}}}}$ denotes the average number of the pair occupation number among five different experiments. As a general trend, the configurations with $\bar{N}_{\scriptstyle{\text{{{{occ}}}}}}=2$ and $N_{\scriptstyle{\text{{{{occ}}}}}}=3$, in that order, are observed. Various factors such as decoherence, cross talk, gate errors and readout errors are expected to contribute to the deviations. In Sec. IV, we will introduce the error mitigation technique modifying the circuits to improve the ratio to obtain the configurations with $N_{\scriptstyle{\text{{{{occ}}}}}}=1$.
@@ V Summary and Outlook
[…]
The “pair-wise” form used in this work has been used in quantum chemistry referred to as the “antisymmetrized geminal power (AGP)” to be used as a reference state of the coupled cluster calculation for the strongly correlated systems to describe two-particle correlation more properly than a single Slater determinant [38, 39, 40, 41, 42, 34, 33, 43], where a geminal is a two-body correspondence of a one-body orbital. The operator of the geminal is quite similar to the Bardeen-Cooper-Schrieffer (BCS) one and, actually, it is pointed out in Ref. [41] that the AGP wave function is strongly related to the particle-number conserving BCS wave function [44]. As an example, Ref. [43] discusses the AGP wave function beyond zero-seniority configurations. Therefore, it is natural to use this AGP ansatz for even-even nuclei as a starting point since the valence nucleons of even-even nuclei form Cooper pairs [45]. It should be noted that the ground-state wave function of the BCS state can also be written by a Pfaffian, instead of as Slater determinant [46, 47, 48]; a Pfaffian wave function and an AGP one are also based on the same philosophy. The number of degrees of freedom involved is different between nuclear physics and quantum chemistry, but the extension of AGP-like ansatz and the techniques developed in quantum chemistry to nuclear many-body problems is an interesting direction to explore.
At the same time, it is interesting to test various hardwares and error mitigation techniques available in the market. The execution time and the decoherence time of NISQ devices are trade-off, i.e., superconducting devices like IBM machines adopted in this work have the faster execution time but the shorter decoherence time, while the trapped-ion devices have the slower execution time but the longer decoherence time. To know the features of each device facilitates for further exploration of the suitable circuits, error mitigation techniques, and the algorithms for the systems of interest.
@@ Acknowledgements.
We acknowledge the use of IBM Quantum services for this work. We thank Dr. Enrico Rinaldi and Prof. Osamu Sugino for the fruitful discussion. This research was supported in part by JSPS Grant-in-Aid for Scientific Research (Grant Nos. JP19H00869, JP20H05670, JP21K18903, JP22H05025, JP22K14030, JP22K20372, JP23H01845, JP23H04526, JP23K01845, JP23K03426, JP24K17057), JST COI-NEXT (Grant No. JPMJPF2221), and MEXT Q-LEAP (Grant No. JPMXS0118067246). S.Y. acknowledges JGC-Saneyoshi Scholarship Foundation. T.N. acknowledges the RIKEN Special Postdoctoral Researcher Program.